OpenStudy (anonymous):

A pizza place offers 5 different cheeses and 10 different toppings. In how many ways can a pizza be made with 2 cheese and 3 toppings? i cannot figure this out and i know it's so simple!

OpenStudy (cwrw238):

do you know how to work out the number of different combinations of 2 from 5?

OpenStudy (anonymous):

5*4*3*2*1?

OpenStudy (anonymous):

i keep getting 120 for this problem

OpenStudy (cwrw238):

no thats factorial 5 or 5!

OpenStudy (anonymous):

5!

OpenStudy (cwrw238):

you need 5C2 = 5! ----- 2! (5-2)!

OpenStudy (anonymous):

i have 5C2 and 10C3

OpenStudy (anonymous):

but i cant figure it out from there

OpenStudy (anonymous):

and i got 10 for 5C2

OpenStudy (cwrw238):

5C2 = 5*4*3*2*1 ---------- 3*2*1 * 2*1

OpenStudy (cwrw238):

a quicker way to do these is as follows 5C2 = 5*4 --- = 10 2*1 10C3 = 10*9*8 ------ 3*2*1

OpenStudy (anonymous):

so 120 times 10?

OpenStudy (cwrw238):

right!

OpenStudy (anonymous):

ok! thanks so much!(:

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

you multiply the 2 results together because the choice of cheeses and toppings are independent events